# A new Angle on Benchmarking Noncovalent Interactions

### v1

@@ 3.8 Assessment of the results when extrapolating to larger molecules
Extrapolating this difference again to the buckycatcher, this would amount to a deviation of 5.6 kJ/mol (and 2.3 kJ/mol for the coronene dimer). Unfortunately, we have CCSDT(Q) values using only a double-zeta basis set. However, we can compute other approximate methods using a larger basis set, compare, and then scale the CCSDT(Q)-CCSD(T) differences in the slopes by these values. For example, the MP2-CCSD(T) slopes differ by 20% when going from a double-zeta basis set to an extrapolated value of triple/quadruple zeta quality. In contrast, the CCSD-CCSD(T) slopes differ by 40% for the same basis sets (DZ and TQZ). As we can see from Table 2, the CCSDT-2 slopes are rather close to CCSDT(Q), and the CCSDT-2-CCSD(T) slopes behave slightly more like CCSD for a small basis set of triple-zeta quality, see Supporting Information. Because of this, we assume that at the double-zeta basis set for the CCSD(T)-CCSDT(Q) slope the difference is underestimated by a maximum of 40%. Hence, we would obtain a deviation of 7.9 kJ/mol for the buckycatcher and 3.2 kJ/mol for the coronene dimer.
@@
For the L7 systems and the buckycatcher dimer in question, the ethylene stacks are not as much of interest, as their post-CCSD(T) effects can be considered as negligible here, the CCSD(T) and CCSDT(Q) slopes are extremely similar with XYZ, arriving at the CCSDT(Q) line y = (-8.7 0.2 kJ/mol)x + 6.5 0.2 kJ/mol per ethylene double bond. In comparison, the final best CCSDT(Q) lines of obtained for the acenes are = (-10.5 0.2 kJ/mol) + 5.2 0.3 kJ/mol , which can also serve as reference for further investigations.
@@ 4 Conclusions
In case of the acene dimer series, CCSD(T) exhibits an increased interaction energy of 3.4% in comparison to CCSDT(Q), which is sizeable, but nowhere near the level that would explain the discrepancies to FN-DMC. For the ethylene stacks, on the other hand, there is no notable difference between CCSD(T) and CCSDT(Q). Thus, with some limitations, we can reconfirm the ”gold standard” of basis set limit CCSD(T) for intermolecular interactions.

### v2

@@ 3.8 Assessment of the results when extrapolating to larger molecules
Extrapolating this difference again to the buckycatcher, this would amount to a deviation of 5.6 kJ/mol (and 2.3 kJ/mol for the coronene dimer). Unfortunately, we have CCSDT(Q) values using only a double-zeta basis set. However, we can compute other approximate methods using a larger basis set, compare, and then scale the CCSDT(Q)-CCSD(T) differences in the slopes by these values. For the acenes, the MP2-CCSD(T) slopes differ by 21% when going from a double-zeta basis set to an extrapolated cc-pV{T,Q}Z value . In contrast, the CCSD-CCSD(T) slopes differ by 57% for the same basis sets (DZ and cc-pV{T,Q}Z). As we can see from Table 2, the CCSDT-2 slopes are rather close to CCSDT(Q), and the CCSDT-2-CCSD(T) slopes behave slightly more like CCSD-CCSD(T) for a small basis set of triple-zeta quality, see Supporting Information. Because of this, we assume that at the double-zeta basis set for the CCSD(T)-CCSDT(Q) slope the difference is underestimated by a maximum of 40%. Hence, we would obtain a deviation from CCSD(T) of 7.9 kJ/mol for the buckycatcher and 3.2 kJ/mol for the coronene dimer.
@@
For the L7 systems and the buckycatcher dimer in question, the ethylene stacks are not as much of interest, as their post-CCSD(T) effects can be considered as negligible here, the CCSD(T) and CCSDT(Q) slopes are extremely similar, arriving at the CCSDT(Q) line y = (-8.7 0.2 kJ/mol)x + 6.5 0.2 kJ/mol per ethylene double bond. In comparison, the final best CCSDT(Q) lines obtained for the acenes are = (-10.5 0.2 kJ/mol) + 5.2 0.3 kJ/mol for the full set of the six acenes. The interaction energies of benzene () and naphthalene () sandwich structures are -6.5 0.2 kJ/mol and -15.5 0.2 kJ/mol, respectively: the resulting slope would be slightly different from the one discussed above because the two data points include the Hartree-Fock association energy, which has its own slope. The interaction energies of ethylene and trans-butadiene, within the geometric constraints applied, are -1.7 kJ/mol and -7.3 kJ/mol. These small molecules or the line of the series can serve as reference for other methods when testing the increase in interaction energy for such species.
@@ 4 Conclusions
In case of the acene dimer series, CCSD(T) exhibits an increased interaction energy of 3.4% in comparison to CCSDT(Q), which is sizeable, but nowhere near the level that would explain the discrepancies to FN-DMC. For the ethylene stacks, on the other hand, there is no notable difference between CCSD(T) and CCSDT(Q). For the ethylene stacks, on the other hand, there is no notable difference between CCSD(T) and CCSDT(Q). Thus, with some limitations, we can reconfirm the ”gold standard” of basis set limit CCSD(T) for intermolecular interactions.